# Interaction-induced phase transitions and critical phases in nonreciprocal non-Hermitian quasicrystals

### v2

@@ Abstract
Non-Hermitian phenomena, such as exceptional points, non-Hermitian skin effects, and topologically nontrivial phases have attracted continued attention. In this work, we reveal how interactions and nonreciprocal hopping could collectively influence the behavior of two interacting bosons on quasiperiodic lattices. Focusing on the Bose-Hubbard model with Aubry-André-Harper quasiperiodic modulations and hopping asymmetry, we discover that interactions could enlarge the localization transition point of the noninteracting system into a critical phase, in which localized doublons formed by bosonic pairs can coexist with delocalized states. Under the open boundary condition, the bosonic doublons could further show non-Hermitian skin effects, realizing doublon condensation at the edges, and their direction of skin-localization can be flexibly tuned by the hopping parameters. A framework is developed to characterize the spectral, localization, and topological transitions accompanying these phenomena. Our work advances the understanding of localization and topological phases in non-Hermitian systems, particularly in relation to multiparticle interactions.
@@ I Introduction
Non-Hermitian systems, usually characterized by non-Hermitian Hamiltonians constitute ideal setups to explore the interplay among localization, environmental effects and interactions. In recent years, non-Hermitian systems have attracted increasing attention due to their distinctive physics, such as the parity-time-reversal (PT) transitions [20, 21, 22], exceptional points [23, 24, 25, 26, 27, 28], non-Hermitian skin effects (NHSEs) [29, 30, 31, 32, 33, 34] and enriched symmetry classifications of topological phases [35, 36, 37, 38, 39, 40, 41]. Meanwhile, the interplay between disorder and non-Hermitian effects may lead to unique types of spectral, localization, topological and entanglement transitions [42, 43, 44, 45, 46, 47, 48, 49, 50, 51, 52, 53, 54, 55, 56, 57, 58, 59, 60, 61, 62, 63, 64, 65, 66, 67, 68, 69, 70, 71, 72, 73, 74, 75], yielding universality classes of localization beyond closed-system contexts [76, 77, 78, 79]. With additional elements like lattice dimerization, long-range hopping, non-Abelian potential and time-period driving, critical phases and mobility edges could also emerge in disordered non-Hermitian systems even in one spatial dimension [80, 81, 82, 83, 84, 85, 86, 87, 88, 89, 90, 91, 92, 93, 94, 95, 96, 97], enriching further the non-Hermitian localization phenomenon.
@@
Interaction could greatly complicates the AL in non-Hermitian systems due to its active competition with other localization mechanisms, including random or correlated disorder and NHSEs. Great efforts have been made towards the understanding of “many-body localization” in non-Hermitian settings, with a focus on multiple interacting particles in lattices of relatively small sizes [98, 99, 100, 101, 102, 103, 104, 105, 106, 107, 108]. Nevertheless, considering a few interacting particles in relatively large lattices may offer necessary insights regarding how interactions could drive different phase transitions and criticality in the thermodynamic limit (with system size $L\rightarrow\infty$) [109, 110, 111, 112, 113, 114, 115, 116, 117, 118]. Motivated by these considerations, we explore in this work the interaction-driven phase transitions and critical phases in non-Hermitian quasicrystals (NHQCs). Focusing on two interacting bosons in one-dimensional (1D) Aubry-André-Harper (AAH) lattices with quasiperiodic onsite potentials and nonreciprocal hoppings, we reveal that the interactions could not only transform the AL critical points of noninteracting systems into critical phases, but also push the boundaries of fully localized phases towards stronger asymmetric hopping regions. Within a critical phase, extended two-boson states can coexist with bounded boson pairs forming doublons, which could also subject to NHSEs under the open boundary condition (OBC). The borders between extended, critical and localized phases are further found to carry topologically nontrivial signatures.
@@
The rest of the paper is organized as follows. In Sec. II, we introduce the models considered in this work, summarizing their known features, and presenting the methods we take to study their physical properties. In Sec. III, we characterize the real-complex spectra transitions, localization transitions, topological transitions and critical phases in representative NHQCs with asymmetric hoppings, tracing their origins back to the interplay between interactions and non-Hermitian effects. In Sec. IV, we focus on the strong-interaction regime and explore the competition of two different localization mechanisms (disorder vs NHSEs) on bosonic doublons. In Sec. V, we summarize our results and discuss potential future directions.
@@ II.1 Model
Figure 1: A sketch of the model [in (a)] and its free-particle phases described by different variables [in (b)–(d), with $U=0$] under the PBC. In (a), we have $V_{l}\equiv 2V\cos(2\pi\alpha l)$, which is quasiperiodic in $l$ for any irrational $\alpha$. (b) shows the ratio of eigenenergies of $\hat{H}_{0}$ with nonzero imaginary parts $\rho_{{\rm Im}}$ [Eq. (3)] vs $(V,J_{{\rm L}})$. (c) shows the averaged IPRs [Eq. (5)] over all eigenstates of $\hat{H}_{0}$ vs $(V,J_{{\rm L}})$. (b) and (c) share the same color bar. (d) gives the topological phase diagram characterized by the winding number $\omega$ [Eq. (6)]. In (b)–(d), the solid line denotes the boundary [Eq. (2)] of PT, localization and topological transitions under the PBC, while the dotted line represents the boundary ($J_{{\rm L}}=J_{{\rm R}}$ and $J_{{\rm L}},J_{{\rm R}}<V$) between left and right skin-localizations of bulk states under the OBC. The hopping amplitude $J_{{\rm R}}=1$ and the length of lattice $L=89$ are set to be the same for (b)–(d).
@@
In the limit of large system size, we have ${\rm IPR}_{{\rm ave}}\rightarrow 0$ if all the eigenstates $|\psi_{j}\rangle$ are spatially extended. On the other hand, if a significant potion $(\propto D)$ of the eigenstates are localized, ${\rm IPR}_{{\rm ave}}$ would take a finite value. In Fig. 1(c), we find a clear boundary between the regions with vanishing and finite ${\rm IPR}_{{\rm ave}}$, which is given by Eq. (2) for the noninteracting $\hat{H}_{0}$. The same boundary can be identified by evaluating the maximum and minimum of IPRs, which confirms that Eq. (2) indeed describes the critical line of localization transition in the quasiperiodic nonreciprocal AAH model. Finally, the winding number $\omega$ counts the number of times that the spectrum of the system winds around a base point $E_{0}^{{\rm B}}$ on the complex energy plane. For the noninteracting $\hat{H}_{0}$, it can be expressed as
@@
To sum up, nonreciprocal non-Hermitian effects could extend the localization transition point of quasiperiodic AAH model [at $J_{\rm L}=V=1$ in Figs. 1(b)–1(d)] into critical lines, through which the system could undergo PT, localization and topological triple phase transitions under the PBC or reverse its skin-localization direction under the OBC. With interactions, the critical point could further expand into a whole critical phase, showing both localized and extended states under the PBC and skin-localized doublons under the OBC, as will be shown in the following section. Before that, we introduce our tools to characterize the phases and transitions of two interacting bosons.
@@ II.2 Methods
To characterize the localization nature of states in real space, we define the symmetrized position eigenbasis as the union of sets $\{|l,l\rangle\}$ and $\{(|l,l^{\prime}\rangle+|l^{\prime},l\rangle)/\sqrt{2}\}$, where $l,l^{\prime}=1,2,...,L$ and $l\neq l^{\prime}$. The first and second indices of each ket denote the positions of two bosons in the lattice. The IPR and its average in Eqs. (4) and (5) can then be computed with respect to this basis. To figure out the existence of critical phases with both extended and localized states, we further introduce the maximum and minimum of IPRs as
@@
In the limit of large system size $L\rightarrow\infty$, we have both ${\rm IPR}_{\max}\rightarrow 0$ and ${\rm IPR}_{\min}\rightarrow 0$ in the extended phase, ${\rm IPR}_{\max}>0$ and ${\rm IPR}_{\min}>0$ in the localized phase, and only ${\rm IPR}_{\min}\rightarrow 0$ in the critical phase. Collecting the information provided by ${\rm IPR}_{\max}$ and ${\rm IPR}_{\min}$ then allows us to distinguish the extended, critical and localized phases. We may also use a single quantity $\zeta$ to identify the presence of critical phases, which is given by [120]
@@
where ${\rm NPR}_{{\rm ave}}\equiv\sum_{j=1}^{D}{\rm IPR}_{j}^{-1}/(LD)$ is the average of normalized participation ratios. In the limit $L\rightarrow\infty$, the function $\zeta$ takes finite values in the critical phase while approaching $\infty$- ∞ in both extended and localized phases. We could then employ ${\rm IPR}_{\max}$, ${\rm IPR}_{\min}$ and $\zeta$ to fully characterize the localization nature of different phases and localization transitions in our system. Note in passing that for better illustrations, our presented values of $\zeta$ are shifted and rescaled as $(\zeta-\min(\zeta))/\max(\zeta-\min(\zeta))$, where the $\min(x)$ and $\max(x)$ find the minimum and maximum of $x$ over the considered domain of system parameters in Figs. 3 and 6 of the next section.
@@
To identify the topological signatures of spectral and localization transitions with two interacting bosons, we note that the single winding number in Eq. (6) may not be enough, as the extended-to-critical and critical-to-localized transitions may possess different phase boundaries. To resolve this issue, we introduce a pair of winding numbers $(\omega_{1},\omega_{2})$, defined versus two different base energies $(E_{1}^{{\rm B}},E_{2}^{{\rm B}})$ as
@@
where the parametrization of $\hat{H}(\phi/L)$ from $\hat{H}$ is the same as that of $\hat{H}_{0}(\phi/L)$ [see the discussion below Eq. (6)]. In numerical calculations, we choose $E_{1}^{{\rm B}}=0$ and set $E_{2}^{{\rm B}}$ as the real part of energy of the eigenstate of $\hat{H}$ with the largest IPR. The quantized change of $\omega_{1}$ from zero to a finite integer then signifies the transition of the spectrum from real to complex. Meanwhile, the quantized jump of $\omega_{2}$ from zero to a finite integer corresponds to the delocalization transition, after which all the eigenstates become extended. Therefore, we would have $\omega_{1}=\omega_{2}=0$ in a phase with real spectrum and only localized states, $\omega_{1},\omega_{2}\neq 0$ in a phase with complex spectrum and only extended states, and $\omega_{1}\neq 0$, $\omega_{2}=0$ in a critical phase with both extended and localized eigenstates. Topological signals of the phase transitions could then be captured by the changes of $\omega_{1}$ or $\omega_{2}$ between zero and finite values.
In the next section, we reveal and characterize the interaction induced spectral transitions, localization transitions, and critical phases in our system from an integrated perspective of eigenspectrum, state distributions and topology, with the help of the tools introduced in this section.
@@ III.1 Unidirectional hopping Model (UHM)
When $U\neq 0$, the spectra in Eq. (12) are modified by the interaction between two bosons. The eigensystem of $\hat{H}_{1}$ can then be obtained numerically by exact diagonalization. In Fig. 2, we present the spectra of $\hat{H}_{1}$ and the IPRs of the associated eigenstates for some typical cases. We find that for $V<J$, a majority of the eigenenergies are complex and confined within the boundary described by Eq. (13). Meanwhile, with the increase of $U$, some eigenvalues are first deformed away from the main part of spectrum, forming a small loop around the highest energy of the noninteracting system, and then collapsed onto the ${\rm Re}E$ axis. Interestingly, the states with real energies generated by a nonzero $U$ are all spatially localized. We thus obtain an interaction-driven critical phase, in which extended and localized eigenstates coexist, which is absent in the noninteracting UHM. The further increase of $U$ will generate at most $L$ localized eigenstates with real energies. When $U$ is large enough, their spectrum would finally be gapped from the extended states with complex energies inside $E_{{\rm b}}$. Each eigenstate in the large $U$ regime is formed by a bosonic doublon localized around one lattice site. When $V>J$, the spectrum of $\hat{H}_{1}$ retains to be real and all its eigenstates are localized for any nonzero $U$. This is not hard to expect, as the onsite Bose-Hubbard interaction could only enhance localization, and the localized particles mainly feel the Hermitian quasiperiodic onsite potential $2V\cos(2\pi\alpha l)$. At large $U$, the spectrum splits into two parts, which are separated by a gap along the ${\rm Re}E$ axis. The collection of states with higher energies are again doublons, with each of them containing two bosons localized on the same lattice site.
@@
Figure 3: Spectral and localization transitions of the UHM vs $(V,U)$ under the PBC. The hopping amplitude is $J=1$ and the length of lattice is $L=89$ for all panels. (a) Ratio of eigenenergies with nonzero imaginary parts $\rho_{{\rm Im}}$ [Eq. (3)]. (b) The maximum of IPRs ${\rm IPR}_{\max}$ [Eq. (7)]. (c) The minimum of IPRs ${\rm IPR}_{\min}$ [Eq. (8)]. (d) The smoking-gun function $\zeta$ [Eq. (9)] of the critical phase.
@@
The above results imply that interactions could transform part of the extended phase of the free-boson UHM into a critical one, which contains both extended two-boson states and localized doublons when the interaction is strong enough. To further confirm this, we present the characteristic functions $\rho_{{\rm Im}}$ [Eq. (3)], ${\rm IPR}_{\max}$ [Eq. (7)], ${\rm IPR}_{\min}$ [Eq. (8)] and $\zeta$ [Eq. (9)] of spectra and states over a range of onsite potential and interaction strengths, as shown in Fig. 3. In Fig. 3(a), we observe a complex-to-real spectral transition at $V=J=1$. The same boundary is found in Fig. 3(c), across which all eigenstates become localized. These observations suggest that the final accomplishment of complex-to-real spectral transition and localization transition are not affected by the onsite Bose-Hubbard interaction in our UHM. However, we notice in Fig. 3(b) a boundary between the regions with ${\rm IPR}_{\max}\simeq 0$ and ${\rm IPR}_{\max}>0$, which is different from the boundary $V=J$ of ${\rm IPR}_{\min}$ in Fig. 3(c) while coinciding with the boundary between $\rho_{{\rm Im}}\simeq 1$ and $10<\rho_{{\rm Im}}<$ 1 in Fig. 3(a). This new phase boundary corresponds to the onset of spectral and localization transitions, which emerges only when $U\neq 0$ and thus has an interaction-origin. However, it could not transform all the states from extended to localized, and the whole spectrum from complex to real. In the region between $V=J$ and the boundary in Fig. 3(b), we now have a critical phase with coexisting localized states of real energies and extended states of complex energies. Such a phase is absent without interactions, and its appearance is further confirmed by the function $\zeta$ in Fig. 3(d), which shows positive values only in the critical phase. Putting together, we confirm that interactions in the UHM $\hat{H}_{1}$ could indeed generate a critical phase and drive the boundary of extended phase towards the region with weaker disorder strengths.
@@
To decode the topological nature of the observed spectral and localization transitions, we compute the winding numbers $(\omega_{1},\omega_{2})$ introduced in Eq. (10), yielding the phase diagram of the UHM in Fig. 4. Within the localized phase, the spectrum of $\hat{H}_{1}$ is real and we find $\omega_{1}=\omega_{2}=0$, as expected. In the critical phase, we obtain $\omega_{1}<0$ and $\omega_{2}=0$. The value of $\omega_{1}$ then undergoes a quantized jump through the real-to-complex spectral transition at $V=J$. In the extended phase, we find both $\omega_{1}<0$ and $\omega_{2}<0$. The value of $\omega_{2}$ thus experiences a quantized change at the boundary through which all the eigenstates become extended. Therefore, the locations where $\omega_{1}$ and $\omega_{2}$ start to deviate from zero in the parameter space could correctly capture all the phase boundaries, offering topological signatures of phase transitions in our system. Note in passing that the exact integer values of $\omega_{1}$ and $\omega_{2}$ rely on the times that the corresponding base energies $E^{\rm B}_{1}$ and $E^{\rm B}_{2}$ are encircled [see Eq. (10)], which could be further dependent on the system size when $\omega_{1}\neq 0$ and/or $\omega_{2}\neq 0$. For the lattice size $L=89$ used in our calculations, we find $\omega_{2}=-2$ ($=0$) in the extended (critical) phase and $\omega_{1}=-27$ in the extended/critical phase, respectively. Therefore, the winding numbers $(\omega_{1},\omega_{2})$ are expected to provide topological signatures of phase transitions in our system, rather than topological characterizations of the phases with different spectral and localization properties.
To sum up, we find that in the UHM, the interaction between two bosons could enlarge the critical point of the noninteracting system at $V=J$ to a critical phase, possessing both extended and localized two-boson states with complex and real energies, respectively. The transition from localized to critical phases also accompanies a real-to-complex spectral transition, while the transition to extended phase is moved up to the domain with weaker disorder in the presence of interactions. In Sec. III.2, we consider the impact of interactions on the more general AAH NHQC with asymmetric hopping amplitudes.

### v3

# Two-body interaction induced phase transitions and intermediate phases in nonreciprocal non-Hermitian quasicrystals

@@ Abstract
Non-Hermitian phenomena, such as exceptional points, non-Hermitian skin effects, and topologically nontrivial phases have attracted continued attention. In this work, we reveal how interactions and nonreciprocal hopping could collectively influence the behavior of two interacting bosons on quasiperiodic lattices. Focusing on the Bose-Hubbard model with Aubry-André-Harper quasiperiodic modulations and hopping asymmetry, we discover that interactions could enlarge the localization transition point of the noninteracting system into an intermediate mobility edge phase, in which localized doublons formed by bosonic pairs can coexist with delocalized states. Under the open boundary condition, the bosonic doublons could further show non-Hermitian skin effects, realizing doublon condensation at the edges, and their direction of skin-localization can be flexibly tuned by the hopping parameters. A framework is developed to characterize the spectral, localization, and topological transitions accompanying these phenomena. Our work advances the understanding of localization and topological phases in non-Hermitian systems, particularly in relation to multiparticle interactions.
@@ I Introduction
Non-Hermitian systems, usually characterized by non-Hermitian Hamiltonians constitute ideal setups to explore the interplay among localization, environmental effects and interactions. In recent years, non-Hermitian systems have attracted increasing attention due to their distinctive physics, such as the parity-time-reversal (PT) transitions [20, 21, 22], exceptional points [23, 24, 25, 26, 27, 28], non-Hermitian skin effects (NHSEs) [29, 30, 31, 32, 33] and enriched symmetry classifications of topological phases [34, 35, 36, 37, 38, 39, 40]. Meanwhile, the interplay between disorder and non-Hermitian effects may lead to unique types of spectral, localization, topological and entanglement transitions [41, 42, 43, 44, 45, 46, 47, 48, 49, 50, 51, 52, 53, 54, 55, 56, 57, 58, 59, 60, 61, 62, 63, 64, 65, 66, 67, 68, 69, 70, 71, 72, 73, 74], yielding universality classes of localization beyond closed-system contexts [75, 76, 77, 78]. With additional elements like lattice dimerization, long-range hopping, non-Abelian potential and time-period driving, intermediate phases and mobility edges could also emerge in disordered non-Hermitian systems even in one spatial dimension [79, 80, 81, 82, 83, 84, 85, 86, 87, 88, 89, 90, 91, 92, 93, 94, 95, 96], enriching further the non-Hermitian localization phenomenon.
@@
Interaction could greatly complicate the AL in non-Hermitian systems due to its active competition with other localization mechanisms, including random or correlated disorder and NHSEs. Great efforts have been made towards the understanding of “many-body localization” in non-Hermitian settings, with a focus on multiple interacting particles in lattices of relatively small sizes [97, 98, 99, 100, 101, 102, 103, 104, 105, 106, 107]. Nevertheless, considering a few interacting particles in relatively large lattices may offer necessary insights regarding how interactions could drive different phase transitions and criticality in the thermodynamic limit (with system size $L\rightarrow\infty$) [108, 109, 110, 111, 112, 113, 114, 115, 116]. Motivated by these considerations, we explore in this work the interaction-driven phase transitions and intermediate mobility edge phases in non-Hermitian quasicrystals (NHQCs). Focusing on two interacting bosons in one-dimensional (1D) Aubry-André-Harper (AAH) lattices with quasiperiodic onsite potentials and nonreciprocal hoppings, we reveal that the interactions could not only transform the AL critical points of noninteracting systems into intermediate mobility edge phases with coexisting extended and localized states, but also push the boundaries of fully localized phases towards stronger asymmetric hopping regions. Within an intermediate phase, extended two-boson states can coexist with bounded boson pairs forming doublons, which could also subject to NHSEs under the open boundary condition (OBC). The borders between extended, intermediate and localized phases are further found to carry topologically nontrivial signatures.
@@
The rest of the paper is organized as follows. In Sec. II, we introduce the model considered in this work, summarizing its known features in the noninteracting limit, and presenting the methods we take to study its physical properties. In Sec. III, we characterize the real-complex spectra transitions, localization transitions, topological transitions and intermediate mobility edge phases in representative NHQCs with asymmetric hoppings, tracing their origins back to the interplay between interactions and non-Hermitian effects of two interacting bosons. In Sec. IV, we focus on the strong-interaction regime and explore the competition of two different localization mechanisms (disorder vs NHSEs) on bosonic doublons. In Sec. V, we summarize our results and discuss potential future directions.
@@ II.1 Model
Figure 1: A sketch of the model [in (a)] and its single-particle phases described by different variables [in (b)–(d), with $U=0$] under the PBC. In (a), we have $V_{l}\equiv 2V\cos(2\pi\alpha l)$, which is quasiperiodic in $l$ for any irrational $\alpha$. (b) shows the ratio of eigenenergies of $\hat{H}_{0}$ with nonzero imaginary parts $\rho_{{\rm Im}}$ [Eq. (3)] vs $(V,J_{{\rm L}})$. (c) shows the averaged IPRs [Eq. (5)] over all eigenstates of $\hat{H}_{0}$ vs $(V,J_{{\rm L}})$. (b) and (c) share the same color bar. (d) gives the topological phase diagram characterized by the winding number $\omega$ [Eq. (6)]. In (b)–(d), the solid line denotes the boundary [Eq. (2)] of PT, localization and topological transitions under the PBC, while the dotted line represents the boundary ($J_{{\rm L}}=J_{{\rm R}}$ and $J_{{\rm L}},J_{{\rm R}}<V$) between left and right skin-localizations of bulk states under the OBC. The hopping amplitude $J_{{\rm R}}=1$ and the length of lattice $L=89$ are set to be the same for (b)–(d).
@@
Note in passing that we only use right eigenvectors and their standard normalization convention in the calculation of IPR. The IPR of each eigenstate then takes values in the range $[0,1]$. In the limit of large system size $L\rightarrow\infty$, we have ${\rm IPR}_{{\rm ave}}\rightarrow 0$ if all the eigenstates $|\psi_{j}\rangle$ are spatially extended, since the IPR of each state goes to zero as $1/L$ in this case. On the other hand, if a significant potion $(\propto D)$ of the eigenstates are localized, the ${\rm IPR}_{{\rm ave}}$ would take a finite value, as the IPR of each localized state is finite in the limit $L\rightarrow\infty$. In Fig. 1(c), we find a clear boundary between the regions with vanishing and finite ${\rm IPR}_{{\rm ave}}$, which is given by Eq. (2) for the noninteracting $\hat{H}_{0}$. The same boundary can be identified by evaluating the maximum and minimum of IPRs, which confirms that Eq. (2) indeed describes the critical line of localization transition in the quasiperiodic nonreciprocal AAH model. Finally, the winding number $\omega$ counts the number of times that the spectrum of the system winds around a base point $E_{0}^{{\rm B}}$ on the complex energy plane. For the noninteracting $\hat{H}_{0}$, it can be expressed as
@@
To sum up, nonreciprocal non-Hermitian effects could extend the localization transition point of quasiperiodic AAH model [at $J_{\rm L}=V=1$ in Figs. 1(b)–1(d)] into critical lines, through which the system could undergo PT, localization and topological triple phase transitions under the PBC or reverse its skin-localization direction under the OBC. With interactions, the critical point could further expand into a whole intermediate phase, showing both localized and extended states separated by mobility edges under the PBC and skin-localized doublons under the OBC, as will be shown in the following section. Before that, we introduce our tools to characterize the phases and transitions of two interacting bosons.
@@ II.2 Methods
In our exact diagonalization calculations, we express the Hamiltonian in Fock space with the occupation number representation $\{|2_{l}\rangle,l=1,...,L\}$ and $\{|1_{l},1_{m}\rangle|l,m=1,...,L;\,l<m\}$. Here, $|2_{l}\rangle$ ($|1_{l},1_{m}\rangle$) denotes the Fock basis with two bosons on the same (different) lattice site $l$ (sites $l$ and $m$). To characterize the localization nature of states in real space, we define the symmetrized position eigenbasis as the union of sets $\{|l,l\rangle\}$ and $\{(|l,m\rangle+|m,l\rangle)/\sqrt{2}\}$, where $l,m=1,2,...,L$ and $l\neq m$. Here, $|l,m\rangle\equiv|l\rangle_{1}\otimes|m\rangle_{2}$ is just the direct product state of position eigenbases in the Hilbert spaces of the two bosons. The first and second indices of each ket $|l,m\rangle$ denote the positions of two bosons in the lattice. The transformations between Fock and position bases are given by $|2_{l}\rangle=|l,l\rangle$ and $|1_{l},1_{m}\rangle=(|l,m\rangle+|m,l\rangle)/\sqrt{2}$. The IPR and its average in Eqs. (4) and (5) can then be computed with respect to the symmetrized position eigenbasis after exact diagonalization. To figure out the existence of intermediate phases with both extended and localized states, we further introduce the maximum and minimum of IPRs as
@@
In the limit of large system size $L\rightarrow\infty$, we have both ${\rm IPR}_{\max}\rightarrow 0$ and ${\rm IPR}_{\min}\rightarrow 0$ in the extended phase, ${\rm IPR}_{\max}>0$ and ${\rm IPR}_{\min}>0$ in the localized phase, and only ${\rm IPR}_{\min}\rightarrow 0$ in the intermediate mobility edge phase. Collecting the information provided by ${\rm IPR}_{\max}$ and ${\rm IPR}_{\min}$ then allows us to distinguish the extended, intermediate and localized phases. We may also use a single quantity $\zeta$ to identify the presence of intermediate mobility edge phases, which is given by [118]
@@
where ${\rm NPR}_{{\rm ave}}\equiv\sum_{j=1}^{D}{\rm IPR}_{j}^{-1}/(LD)$ is the average of normalized participation ratios. In the limit $L\rightarrow\infty$, the function $\zeta$ takes finite values in the intermediate mobility edge phase while approaching $\infty$- ∞ in both extended and localized phases. We could then employ ${\rm IPR}_{\max}$, ${\rm IPR}_{\min}$ and $\zeta$ to fully characterize the localization nature of different phases and localization transitions in our system. Note in passing that for better illustrations, our presented values of $\zeta$ are shifted and rescaled as $(\zeta-\min(\zeta))/\max(\zeta-\min(\zeta))$, where the $\min(x)$ and $\max(x)$ find the minimum and maximum of $x$ over the considered domain of system parameters in Figs. 3 and 6 of the next section.
@@
To identify the topological signatures of spectral and localization transitions with two interacting bosons, we note that the single winding number in Eq. (6) may not be enough, as the extended-to-intermediate and intermediate-to-localized transitions may possess different phase boundaries. To resolve this issue, we introduce a pair of winding numbers $(\omega_{1},\omega_{2})$, defined versus two different base energies $(E_{1}^{{\rm B}},E_{2}^{{\rm B}})$ as
@@
where the parametrization of $\hat{H}(\phi/L)$ from $\hat{H}$ is the same as that of $\hat{H}_{0}(\phi/L)$ [see the discussion below Eq. (6)]. In numerical calculations, we choose $E_{1}^{{\rm B}}=0$ and set $E_{2}^{{\rm B}}$ as the real part of energy of the eigenstate of $\hat{H}$ with the largest IPR. The quantized change of $\omega_{1}$ from zero to a finite integer then signifies the transition of the spectrum from real to complex. Meanwhile, the quantized jump of $\omega_{2}$ from zero to a finite integer corresponds to the delocalization transition, after which all the eigenstates become extended. Therefore, we would have $\omega_{1}=\omega_{2}=0$ in a phase with real spectrum and only localized states, $\omega_{1},\omega_{2}\neq 0$ in a phase with complex spectrum and only extended states, and $\omega_{1}\neq 0$, $\omega_{2}=0$ in an intermediate mobility edge phase with both extended and localized eigenstates. Topological signals of the phase transitions could then be captured by the changes of $\omega_{1}$ or $\omega_{2}$ between zero and finite values.
In the next section, we reveal and characterize the interaction induced spectral transitions, localization transitions, and intermediate phases in our system from an integrated perspective of eigenspectrum, state distributions and topology, with the help of the tools introduced in this section.
@@ III.1 Unidirectional hopping Model (UHM)
When $U\neq 0$, the spectra in Eq. (12) are modified by the interaction between two bosons. The eigensystem of $\hat{H}_{1}$ can then be obtained numerically by exact diagonalization. In Fig. 2, we present the spectra of $\hat{H}_{1}$ and the IPRs of the associated eigenstates for some typical cases. We find that for $V<J$, a majority of the eigenenergies are complex and confined within the boundary described by Eq. (13). Meanwhile, with the increase of $U$, some eigenvalues are first deformed away from the main part of spectrum, forming a small loop around the highest energy of the noninteracting system, and then collapsed onto the ${\rm Re}E$ axis. Interestingly, the states with real energies generated by a nonzero $U$ are all spatially localized. We thus obtain an interaction-driven intermediate mobility edge phase, in which extended and localized eigenstates coexist, which is absent in the noninteracting UHM. The further increase of $U$ will generate at most $L$ localized eigenstates with real energies. When $U$ is large enough, their spectrum would finally be gapped from the extended states with complex energies inside $E_{{\rm b}}$. Each eigenstate in the large $U$ regime is formed by a bosonic doublon localized around one lattice site. When $V>J$, the spectrum of $\hat{H}_{1}$ retains to be real and all its eigenstates are localized for any nonzero $U$. This is not hard to expect, as the onsite Bose-Hubbard interaction could only enhance localization, and the localized particles mainly feel the Hermitian quasiperiodic onsite potential $2V\cos(2\pi\alpha l)$. At large $U$, the spectrum splits into two parts, which are separated by a gap along the ${\rm Re}E$ axis. The collection of states with higher energies are again doublons, with each of them containing two bosons localized on the same lattice site.
@@
Figure 3: Spectral and localization transitions of the UHM vs $(V,U)$ under the PBC. The hopping amplitude is $J=1$ and the length of lattice is $L=89$ for all panels. (a) Ratio of eigenenergies with nonzero imaginary parts $\rho_{{\rm Im}}$ [Eq. (3)]. (b) The maximum of IPRs ${\rm IPR}_{\max}$ [Eq. (7)]. (c) The minimum of IPRs ${\rm IPR}_{\min}$ [Eq. (8)]. (d) The smoking-gun function $\zeta$ [Eq. (9)] of the intermediate mobility edge phase.
@@
The above results imply that interactions could transform part of the extended phase of the free-boson UHM into a intermediate one, which contains both extended two-boson states and localized doublons when the interaction is strong enough. To further confirm this, we present the characteristic functions $\rho_{{\rm Im}}$ [Eq. (3)], ${\rm IPR}_{\max}$ [Eq. (7)], ${\rm IPR}_{\min}$ [Eq. (8)] and $\zeta$ [Eq. (9)] of spectra and states over a range of onsite potential and interaction strengths, as shown in Fig. 3. In Fig. 3(a), we observe a complex-to-real spectral transition at $V=J=1$. The same boundary is found in Fig. 3(c), across which all eigenstates become localized. These observations suggest that the final accomplishment of complex-to-real spectral transition and localization transition are not affected by the onsite Bose-Hubbard interaction in our UHM. However, we notice in Fig. 3(b) a boundary between the regions with ${\rm IPR}_{\max}\simeq 0$ and ${\rm IPR}_{\max}>0$, which is different from the boundary $V=J$ of ${\rm IPR}_{\min}$ in Fig. 3(c) while coinciding with the boundary between $\rho_{{\rm Im}}\simeq 1$ and $10<\rho_{{\rm Im}}<$ 1 in Fig. 3(a). This new phase boundary corresponds to the onset of spectral and localization transitions, which emerges only when $U\neq 0$ and thus has an interaction-origin. However, it could not transform all the states from extended to localized, and the whole spectrum from complex to real. In the region between $V=J$ and the boundary in Fig. 3(b), we now have an intermediate phase with coexisting localized states of real energies and extended states of complex energies. Such a phase is absent without interactions, and its appearance is further confirmed by the function $\zeta$ in Fig. 3(d), which shows positive values only in the intermediate mobility edge phase. Putting together, we confirm that interactions in the UHM $\hat{H}_{1}$ could indeed generate a mobility edge phase and drive the boundary of extended phase towards the region with weaker disorder strengths.
@@
To decode the topological nature of the observed spectral and localization transitions, we compute the winding numbers $(\omega_{1},\omega_{2})$ introduced in Eq. (10), yielding the phase diagram of the UHM in Fig. 4. Within the localized phase, the spectrum of $\hat{H}_{1}$ is real and we find $\omega_{1}=\omega_{2}=0$, as expected. In the intermediate mobility edge phase, we obtain $\omega_{1}<0$ and $\omega_{2}=0$. The value of $\omega_{1}$ then undergoes a quantized jump through the real-to-complex spectral transition at $V=J$. In the extended phase, we find both $\omega_{1}<0$ and $\omega_{2}<0$. The value of $\omega_{2}$ thus experiences a quantized change at the boundary through which all the eigenstates become extended. Therefore, the locations where $\omega_{1}$ and $\omega_{2}$ start to deviate from zero in the parameter space could correctly capture all the phase boundaries, offering topological signatures of phase transitions in our system. Note in passing that the exact integer values of $\omega_{1}$ and $\omega_{2}$ rely on the times that the corresponding base energies $E^{\rm B}_{1}$ and $E^{\rm B}_{2}$ are encircled [see Eq. (10)], which could be further dependent on the system size when $\omega_{1}\neq 0$ and/or $\omega_{2}\neq 0$. For the lattice size $L=89$ used in our calculations, we find $\omega_{2}=-2$ ($=0$) in the extended (intermediate) phase and $\omega_{1}=-27$ in the extended/intermediate phase, respectively. Therefore, the winding numbers $(\omega_{1},\omega_{2})$ are expected to provide topological signatures of phase transitions in our system, rather than topological characterizations of the phases with different spectral and localization properties.
To sum up, we find that in the UHM, the interaction between two bosons could enlarge the critical point of the noninteracting system at $V=J$ to an intermediate mobility edge phase, possessing both extended and localized two-boson states with complex and real energies, respectively. The transition from localized to intermediate phases also accompanies a real-to-complex spectral transition, while the transition to extended phase is moved up to the domain with weaker disorder in the presence of interactions. In Sec. III.2, we consider the impact of interactions on the more general AAH NHQC with asymmetric hopping amplitudes.